# Testing the cosmological Euler equation: viscosity, equivalence principle, and gravity beyond general relativity

### v1

@@ IV Viscous dark matter and galaxy clustering
The parameter $E_{P}$ has been previously introduced in Castello:2024lhl as a model-independent way to test the equivalence principle. Here, we see that it is generalized to $\tilde{E}_{P}$ when including viscosity. We will frequently use the parameter $C_{\mathrm{vis}}$ as a measure of viscosity in the following and denote its present value at redshift zero with $C_{\mathrm{vis},0}$, a quantity that, as we will see, plays a central role throughout this paper. In Appendix G, we provide upper limits on the magnitude of $C_{\mathrm{vis},0}$, derived from existing observational constraints on dark matter viscosity.
@@
However, as we show in the forecast analysis below, the viscosity parameter $C_{\mathrm{vis},0}$ is expected to be constrained to very small values, implying that both $\mathcal{Z}$ and $\mathcal{E}$ are very small at all epochs. In this case, and also assuming for simplicity that $\eta,\mu_{G}$ are constant, we can rewrite $\tilde{E}_{P}$ as  [^5] Here we neglect terms quadratic in viscosities and where $\mathcal{Z}$ or $\mathcal{E}$ is multiplied by $\Theta$, $\Gamma$ or $1-\eta$.
@@ VII Results and Discussion
We present the forecasting results for the baseline bias configuration with $b_{g,F}/b_{g,B}=0.5$ in Tables 8, 9, and 10, corresponding to the DESI, Euclid, and SKA2 cases, respectively. We begin by commenting on the measurability of the viscosity parameter $C_{\rm vis,0}$. From these tables, we find that upcoming galaxy surveys will be capable of placing robust and remarkably precise constraints on the viscosity of dark matter. Namely, the present-day value of the dimensionless viscosity parameter $C_{\rm vis,0}$ can be constrained at the level of $\mathcal{O}(10^{-7})$, with the exact sensitivity depending on the specific survey configurations. We remark that our constraints are of a comparable order of magnitude to previous constraints (for a more detailed discussion on these constraints, see Appendix G), but with many fewer model assumptions.
@@ Appendix G Observational constraints for viscosity
Normann:2016jns $\tilde{\zeta}\lesssim 2.6\times 10^{-2}$ $C_{vis,0}\lesssim 0.9\times 10^{-2}$ independent $H(z)$ data from Ref. Chen:2013vea $\zeta$ constant; $\zeta(z)$
Velten:2012uv $\tilde{\zeta}\leq 0.31$ ($2\,\sigma$ c.l.) $C_{vis,0}\leq 0.11$ SN, BAO, CMB
$\tilde{\zeta}\leq 10^{-6}$ $C_{vis,0}\leq 3.5\times 10^{-7}$ $\sim$ 10 Mpc clusters $\zeta$ constant
$\tilde{\zeta}\leq 10^{-9}$ $C_{vis,0}\leq 3.5\times 10^{-10}$ $\sim$ 10 kpc galaxies
$\tilde{\zeta}\leq 10^{-11}$ $C_{vis,0}\leq 3.5\times 10^{-12}$ $\sim$ 1 kpc dwarf galaxies
Velten:2013pra $\tilde{\zeta}\leq 5\times 10^{-11}$ $C_{vis,0}\leq 1.8\times 10^{-11}$ $\sim$ 1 kpc dwarf galaxies $\zeta$ constant; $\zeta(z)$; (neo-)Newtonian and relativistic approaches
Velten:2014xca $\tilde{\zeta}\approx 10^{-8}$ $C_{vis,0}\approx 10^{-9}$ linear power spectrum and cumulative mass function $\zeta$ constant
Barbosa:2017ojt $\tilde{\zeta}\leq 1.427\times 10^{-6}$ $\tilde{\eta}_{s}\leq 2.593\times 10^{-6}$ $C_{vis,0}\leq 1.7\times 10^{-6}$ $f\sigma_{8}(z)$ data from RSD measurements $\zeta$ and $\eta_{s}$ constant
Ashoorioon:2023jwf $\tilde{\zeta}=0.26^{+0.05}_{-0.04}\times 10^{-6}$ $C_{vis,0}=(0.92\pm 0.16)\times 10^{-7}$ CMB, BAO, SN, CC, LSS $\zeta$ constant; $\zeta(z)$
@@
Anand:2017wsj $\tilde{\zeta}=1.32^{+0.50}_{-1.00}\times 10^{-6}$ $\tilde{\eta}_{s}=1.20^{+0.40}_{-1.00}\times 10^{-6}$ $C_{vis,0}=(1.03\pm 0.45)\times 10^{-6}$ CMB, SZ, lensing, BAO $\zeta$ and $\eta_{s}$ constant
Goswami:2016tsu $\tilde{\eta}_{s}\leq 59.2$ $C_{vis,0}\leq 27.8$ dissipation of gravitational waves from GW150914 GW propagation through fluid with $\eta_{s}$ constant

### v2

@@ IV Viscous dark matter and galaxy clustering
[…]
@@
On the other hand, as we show in the forecast analysis below, the viscosity parameter $C_{\mathrm{vis},0}$ is expected to be constrained to very small values, implying that both $\mathcal{Z}$ and $\mathcal{E}$ are very small at all epochs. In this case, and also assuming for simplicity that $\eta,\mu_{G}$ are constant, we can rewrite $\tilde{E}_{P}$ as  [^5] Here we neglect terms quadratic in viscosities and where $\mathcal{Z}$ or $\mathcal{E}$ is multiplied by $\Theta$, $\Gamma$ or $1-\eta$.
@@ VII Results and Discussion
We present the forecasting results for the baseline bias configuration with $b_{g,F}/b_{g,B}=0.5$ in Tables 8, 9, and 10, corresponding to the DESI, Euclid, and SKA2 cases, respectively. All results displayed below are obtained by marginalizing over all the remaining parameters. We begin by commenting on the measurability of the viscosity parameter $C_{\rm vis,0}$. From these tables, we find that upcoming galaxy surveys will be capable of placing robust and remarkably precise constraints on the viscosity of dark matter. Namely, the present-day value of the dimensionless viscosity parameter $C_{\rm vis,0}$ can be constrained at the level of $\mathcal{O}(10^{-7})$, with the exact sensitivity depending on the specific survey configurations. We remark that our constraints are of a comparable order of magnitude to previous constraints (for a more detailed discussion on these constraints, see Appendix G), but with many fewer model assumptions.
@@ Appendix G Observational constraints for viscosity
Normann:2016jns $\tilde{\zeta}\lesssim 2.6\times 10^{-2}$ $C_{\mathrm{vis},0}\lesssim 0.9\times 10^{-2}$ independent $H(z)$ data from Ref. Chen:2013vea $\zeta$ constant; $\zeta(z)$
Velten:2012uv $\tilde{\zeta}\leq 0.31$ ($2\,\sigma$ c.l.) $C_{\mathrm{vis},0}\leq 0.11$ SN, BAO, CMB
$\tilde{\zeta}\leq 10^{-6}$ $C_{\mathrm{vis},0}\leq 3.5\times 10^{-7}$ $\sim$ 10 Mpc clusters $\zeta$ constant
$\tilde{\zeta}\leq 10^{-9}$ $C_{\mathrm{vis},0}\leq 3.5\times 10^{-10}$ $\sim$ 10 kpc galaxies
$\tilde{\zeta}\leq 10^{-11}$ $C_{\mathrm{vis},0}\leq 3.5\times 10^{-12}$ $\sim$ 1 kpc dwarf galaxies
Velten:2013pra $\tilde{\zeta}\leq 5\times 10^{-11}$ $C_{\mathrm{vis},0}\leq 1.8\times 10^{-11}$ $\sim$ 1 kpc dwarf galaxies $\zeta$ constant; $\zeta(z)$; (neo-)Newtonian and relativistic approaches
Velten:2014xca $\tilde{\zeta}\approx 10^{-8}$ $C_{\mathrm{vis},0}\approx 10^{-9}$ linear power spectrum and cumulative mass function $\zeta$ constant
Barbosa:2017ojt $\tilde{\zeta}\leq 1.427\times 10^{-6}$ $\tilde{\eta}_{s}\leq 2.593\times 10^{-6}$ $C_{\mathrm{vis},0}\leq 1.7\times 10^{-6}$ $f\sigma_{8}(z)$ data from RSD measurements $\zeta$ and $\eta_{s}$ constant
Ashoorioon:2023jwf $\tilde{\zeta}=0.26^{+0.05}_{-0.04}\times 10^{-6}$ $C_{\mathrm{vis},0}=(0.92\pm 0.16)\times 10^{-7}$ CMB, BAO, SN, CC, LSS $\zeta$ constant; $\zeta(z)$
@@
Anand:2017wsj $\tilde{\zeta}=1.32^{+0.50}_{-1.00}\times 10^{-6}$ $\tilde{\eta}_{s}=1.20^{+0.40}_{-1.00}\times 10^{-6}$ $C_{\mathrm{vis},0}=(1.03\pm 0.45)\times 10^{-6}$ CMB, SZ, lensing, BAO $\zeta$ and $\eta_{s}$ constant
Goswami:2016tsu $\tilde{\eta}_{s}\leq 59.2$ $C_{\mathrm{vis},0}\leq 27.8$ dissipation of gravitational waves from GW150914 GW propagation through fluid with $\eta_{s}$ constant